OpenStudy (anonymous):

F(x)=5x^2-3x+6 Find f(k-1)

OpenStudy (jack1):

F(x) = 5x^2 - 3x + 6 (k-1) \[F(x) = 5x^2 - 3x + 6 \] \[F((k-1)) = 5(k-1)^2 - 3(k-1) + 6 \] \[F((k-1)) = 5(k-1)(k-1) - 3(k-1) + 6 \] \[F((k-1)) = 5(k-1)(k-1) - 3k-3 + 6 \] \[F((k-1)) = 5(k^2 -1k-1k + 1) - 3k-3 + 6 \] \[F((k-1)) = 5(k^2 -2k + 1) - 3k-3 + 6 \] \[F((k-1)) = 5k^2 -10k + 5 - 3k-3 + 6 \] can you solve from here dude @blackhawkeye47 ???